OpenStudy (chrisneedshelp):

Set up a proportion and use cross multiplication to solve. 15 is ?% of 60.

OpenStudy (anonymous):

Let the percentage be : x x % of 60 = 15 \[\frac{x}{100} \times 60 = 15\] \[x = \frac{15 \times 100}{60}\] Can you solve it now??

OpenStudy (chrisneedshelp):

25

OpenStudy (anonymous):

Yes you are right..

OpenStudy (chrisneedshelp):

thank you

OpenStudy (anonymous):

Welcome..